a park ranger caught. tagged and released 612 deer. He then caught 244 deer from the same locaction. 72 of those deer were tagged. How many deer are in that location? The answer is 2074 deer; however, I do not know how to arrive at that answer. What is the constant and how is it derived?
Answer by checkley77(12844)

612/X=72/244
72X=612*244
72X=149,328
X=149,328/72
X=2,074 ANS.